OpenStudy (frostbite):

Are logarithms the only function for which f(xy)=f(x)+f(y) is true? (Mathematical foundations of thermodynamics)

OpenStudy (frostbite):

In addition we require the function to be continuous and non-discrete.

OpenStudy (frostbite):

@Michele_Laino

OpenStudy (michele_laino):

from such identity, comes the additivity of the entropy

OpenStudy (frostbite):

So far I have figured out that yes. I was simply wondering if there were other functions in an attempt to understand why it ended being the natural logarithm that was used. @Michele_Laino

OpenStudy (rational):

The second thing that comes to my mind after logs is the argument of a complex number : \[\arg(z_1z_2) = \arg(z_1)+\arg(z_2)\]

OpenStudy (anonymous):

Also, not too helpful, but the trivial constant functions \(f(x)=0\) and \(f(x)=2\) work too (though is the second one actually trivial?).

OpenStudy (kainui):

There's a number theory function that obeys this relationship but for positive integers only, \[\Omega(a*b)=\Omega(a)+\Omega(b)\] \(\Omega(n)\) is the number of prime divisors in n, for instance, \(\Omega(12)=\Omega(2*2*3) = 3\) since it has 3 prime divisors, or we can expand it like this: \(\Omega(12)=\Omega(2*2*3)=\Omega(2)+\Omega(2)+\Omega(3) = 1+1+1\) Although it's not so much different than the logarithm function, really.

OpenStudy (thomas5267):

https://en.wikipedia.org/wiki/Logarithm#Logarithmic_function More precisely, the logarithm to any base b > 1 is the only increasing function f from the positive reals to the reals satisfying f(b) = 1 and [31]\[ f(xy)=f(x)+f(y)\]